OpenStudy (anonymous):

How do you do exponential expressions without using a calculator? 2^-3

OpenStudy (jdoe0001):

\(\bf \large a^{-n} = \cfrac{1}{a^n}\)

OpenStudy (anonymous):

Thank you so much. So 2^-3 would be 1/2^3= 1/6?

OpenStudy (jdoe0001):

sorta, \(\bf 2 \times 3 = 6\qquad \qquad 2^3=8\)

OpenStudy (anonymous):

Oh yeah, duh. What if it was a fraction like \[(1/2)^{-4}\]?

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{1}{4}\right)^{-4}\implies \cfrac{1}{\left(\frac{1}{4}\right)^{4}}\implies \cfrac{\frac{1}{1}}{\frac{1^4}{4^4}}\implies \cfrac{1}{1}\times \cfrac{4^4}{1^4}\)

OpenStudy (jdoe0001):

hmmm shoot.. I used ... 4 dunno why hehhe, typo I gather, lemme fix that

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{1}{2}\right)^{-4}\implies \cfrac{1}{\left(\frac{1}{2}\right)^{4}}\implies \cfrac{\frac{1}{1}}{\frac{1^4}{2^4}}\implies \cfrac{1}{1}\times \cfrac{2^4}{1^4}\)